OpenStudy (anonymous):

Solve: y^3 + 6y^2 = 27y

OpenStudy (anonymous):

y(y^2+6y-27)=0 y(y^2 +9y -3y -27)=0 y { y(y+9) -3(y+9)}=0 y{ (y+9)(y-3)}=0 y(y+9)(y-3)=0 hence y=0,-9,3

OpenStudy (anonymous):

y^3 +6y^2-27y=0 y(y-3)(y+9)=0 y=-9 y=0 y=3

OpenStudy (anonymous):

=> y^3 + 6y^2 - 27Y = 0 => y(Y^2 + 6y - 27) = 0 => y [ y^2 + 9y -3y - 27) = 0 => y[ y(y+9) -3(y+9) = 0 => y (y-3) (y+9) = 0

OpenStudy (anonymous):

the product can be zero if either of the factors r zero y = 0 => y=0 y-3 = 0 => y = 3 y+9 = 0 => y = -9 so y = 0, 3, -9